3 1/3 divided by 1 5/9
Answer by richard1234(7193)

Convert to improper fraction, i.e. 10/3 divided by 14/9. To divide, we invert the divisor (14/9) and multiply.

(10/3)(9/14) = 90/42 = 15/7.